# On the Adversarial Risk of Test Time Adaptation: An Investigation into Realistic Test-Time Data Poisoning

### v2

@@ 3.3 TTA-Aware Attack Objective
Low Entropy Attack Objective: High entropy attack objective is particular effective against self-training based TTA methods because of high updating gradient, yet they could be easily defended by some defense strategies, e.g. entropy thresholding. Therefore, we further explore a new low entropy based attack objective. DIA (Wu et al., 2023) proposed to maximize the cross-entropy loss on the benign samples (indiscriminate attack) to generate the other poisoned samples. However, we empirically found that maximizing the cross-entropy loss without any constraints on one sample is prone to maximizing the probability of the most confident class (except the ground-truth) of one model, and feeding these samples into TTA model would bring up the following issues. i) The model will quickly bias towards the most confident class and collapse if without any class diversity constraints. ii) If class diversity constraints is applied (assembled in several TTA methods, e.g. EATA, ROID), this objective will become less ineffective since the class-biased poisoned samples will obtain the less updating weighting than other benign samples. Therefore, we propose a class-balanced low entropy attack, termed Balanced Low Entropy Attack (BLE Attack). Specifically, we maintain an moving average probability confusion $C\in[0,1]^{K\times K}$ to store the prediction bias in each class and find a global optimal label mapping $M\in\{0,1\}^{K\times K}$ such that each class is attacked towards the most probable non ground-truth class. Details of deriving label mapping $M$ is deferred to the supplementary. Finally, the BLE Attack objective is calculate as Eq. 8.
@@
Overall Attack Strategy: We craft poisoned data by attacking the aforementioned in-distribution attack objective with regularization. Following the practice that poisoned data is less discernible by human, we employ PGD attack (Madry et al., 2018) on the combined objective in Eq. 6 with a budget $b$. More details of the training and evaluation algorithm are deferred to the Appendix.
@@ 4.1 Experiment Details
Benchmark Poisoning Methods: We evaluated the following methods under our proposed RTTDP setting. DIA (Wu et al., 2023) is one of the first approaches towards test-time data poisoning, generating the poisoned data via maximizing the cross-entropy of other benign data. We adapt DIA to the realistic evaluation protocol by splitting $\mathcal{B}_{ab}$ into two subsets of $50\%$ each, i.e. $\mathcal{B}^{p}_{ab}:\mathcal{B}^{b}_{ab}=1:1$, and craft $\mathcal{B}^{p}_{a}$ to maximize the cross-entropy of $\mathcal{B}^{b}_{ab}$. TePA (Cong et al., 2023) proposed to maximize entropy to generate poisoned data and performed attack before TTA starts. We adapt TePA to generate the poisoned data based on the source model and inject them into TTA pipeline on-the-fly. MaxCE (Goodfellow et al., 2014) is an established way to create adversarial samples by maximizing the cross-entropy loss. Finally, we evaluate the two attack objectives proposed in this paper, i.e high entropy attack (NHE Attack) and low entropy attack against most probable and balanced non ground-truth class (BLE Attack). For both NHE Attack and BLE Attack, we evaluate the attack objective subject to the constraint of our proposed feature consistency (Eq. 6).
Datasets: We evaluate on three datasets, widely adopted for TTA benchmarking. CIFAR10-C, CIFAR100-C and ImageNet-C are synthesized from the original clean validation set by adding various types of corruptions to simulate natural distribution shifts (Hendrycks & Dietterich, 2019). We choose corruption level 5 and perform continual test-time adaptation setting for evaluation. Following prior works (Wang et al., 2022; Döbler et al., 2023), we adopt the pre-trained WideResNet-28 Zagoruyko & Komodakis (2016), ResNeXt-29 (Xie et al., 2017), and ResNet-50 (He et al., 2016) models for experiments on the CIFAR10-C, CIFAR100-C, and ImageNet-C datasets, respectively. More experiment details can be found in the Appendix.
@@ 4.2 Evaluation on Test-Time Data Poisoning
Table 2: Evaluation of test-time data poisoning under the RTTDP protocol for CIFAR10-C. We report the attack success rate (higher the better) for each TTA method and the average ranking (lower the better) for each attack objective.
@@
Uniform No Attack 43.81 19.72 21.00 18.03 17.41 18.94 16.46 16.37 18.28 5.71
DIA (Wu et al., 2023) 26.04 21.87 18.94 18.56 19.46 17.72 17.77 20.05 3.71
TePA (Cong et al., 2023) 33.78 22.36 23.37 17.75 19.53 16.57 18.76 21.73 3.43
MaxCE (Goodfellow et al., 2014) 18.55 20.81 18.17 18.50 19.50 16.88 18.57 18.71 4.57
BLE Attack (Ours) 54.07 51.99 45.20 34.00 26.80 18.12 19.06 35.61 1.29
NHE Attack (Ours) 73.86 72.40 29.73 18.67 24.56 17.54 17.00 36.25 2.29
@@
Table 3: Evaluation of test-time data poisoning under the RTTDP protocol for CIFAR100-C. We report the attack success rate (higher the better) for each TTA method and the average ranking (lower the better) for each attack objective.
@@
Uniform No Attack 46.23 60.25 47.12 32.20 31.93 31.62 32.13 29.11 37.77 5.71
DIA (Wu et al., 2023) 74.79 68.42 33.77 32.57 33.20 32.68 30.10 43.65 4.00
TePA (Cong et al., 2023) 75.79 81.98 36.74 33.41 35.79 32.41 32.24 46.91 2.86
MaxCE (Goodfellow et al., 2014) 42.36 39.04 34.08 40.73 32.01 32.18 31.08 35.93 4.43
BLE Attack (Ours) 73.93 76.71 47.30 34.58 43.25 32.81 32.27 48.69 2.00
NHE Attack (Ours) 92.08 91.72 37.86 33.85 56.09 32.50 31.48 53.65 2.00
@@
Table 4: Evaluation of test-time data poisoning under the RTTDP protocol for ImageNet-C. We report the attack success rate (higher the better) for each TTA method and the average ranking (lower the better) for each attack objective.
@@ 4.3 Ablation Study on Attack Modules
In this section, we ablate our proposed modules including the surrogate model, the feature consistency regularization and two attack objectives to demonstrate their their indispensable contribution to the final results. We conduct the experiments on CIFAR10-C and CIFAR100-C datasets as shown in Tab. 5. First, comparing using source model and surrogate model to generate poisoned data, using surrogate model would give superior results which approach or even surpass slightly the ones using online model, regardless of the attack objective. It demonstrates the effectiveness of our proposed surrogate model that is leveraged for generating on-the-fly poisoned data. Second, our proposed NHE attack objective is effective though using source model and without feature consistency regularization, that could be attributed to the high entropy samples easily mislead the model update and cause strong perturbation to the source knowledge. Third, under the surrogate model or online model, both BLE and NHE are significantly improved with the help of feature consistency regularization, empirically demonstrating the reasonableness and effectiveness of our method.
@@
TENT EATA SAR Avg. Err. ($\uparrow$) TENT EATA SAR Avg. Err. ($\uparrow$)
@@ 4.5 Generalisation of our modules
In this section, to demonstrate the generalisation of our method, we combine the existing attack objectives with our proposed modules including the surrogate model and feature consistency regularization. The comparisons are shown in Tab. 7. First, we can observe that DIA and MaxCE could get improved additional with our proposed method. Second, TePA could obtain improvement under TENT method but slightly degraded under EATA method. It could be that TePA using maximizing entropy as an attack objective, and with the help of the surrogate model, the poisoned data would have very high prediction entropy because the attack reference model is more approximate to the online model, but fail to pass the entropy threshold and class diverse weighting using in EATA. Overall, our proposed in-distribution attack could generalize to most of the existing attacking objectives.
5 Related Work
Test-Time Adaptation (TTA)  (Wang et al., 2020; 2022; Niu et al., 2022; Su et al., 2022; Li et al., 2023) have shown significant success in bridging the domain gap by using stream-based testing samples to dynamically update models in real time. The success of TTA is mainly attributed to the self-supervised learning on testing data. While it has proved sensitive to confirmation bias (Arazo et al., 2020), many solutions were proposed to minimize the influences of the wrong pseudo-labels. These strategies include minimizing sample entropy (Wang et al., 2020; Liang et al., 2020), adding regularization terms (Song et al., 2023; Su et al., 2024b), using confidence thresholding (Niu et al., 2022; 2023), updating models with exponential moving average architectures (Wang et al., 2022; Döbler et al., 2023), partially updating model weights (Wang et al., 2020; Yuan et al., 2023), and augmenting testing samples (Zhang et al., 2022; Döbler et al., 2023), all aimed at minimizing sample distribution discrepancies. However, adaptation during the testing stage remains highly risky, as it directly impacts the final results. In this work, we aim to demonstrate that most existing TTA methods are highly susceptible to data poisoning.
Adversarial Attack & Data Poisoning: Adversarial risk is a crucial concern for models to address for safe deployment, which can be primarily divided into two categories: adversarial attacks and data poisoning. Adversarial attacks (Szegedy et al., 2014; Akhtar & Mian, 2018; Goodfellow et al., 2014; Madry et al., 2018) manipulate models during inference by adding small perturbations to input data. White-box attacks (Szegedy et al., 2014; Tramèr et al., 2018) assume full access to the victim model, creating adversarial examples by maximizing loss gradients. In contrast, grey-box attacks (Chen et al., 2017; Ilyas et al., 2018; Ru et al., 2019) assume no model access, generating adversarial samples by estimating gradients through intensive querying. Data poisoning (Biggio et al., 2012; Yang et al., 2017; Shafahi et al., 2018; Alfeld et al., 2016) compromise models by injecting manipulated data into the training set, misleading the training process. Traditional data poisoning assumes that the attacker can observe and poison the entire training set at once. Recent advancements, such as online poisoning (Zhang et al., 2020), relax this assumption by requiring knowledge of the model updating strategies. In this work, we explore a more realistic scenario of adversarial robustness for TTA, focusing on data poisoning without online model access and without multiple queries for the same data.
@@ 6 Conclusion
Limitation: We list the limitations as follows. The proposed test-time data poisoning strategy is based on grey-box attack, i.e. model architecture and initial weights are available to the adversary. The proposed attack does not guarantee huge improvement of attack success rate on all TTA methods, especially the ones integrated with multiple defense techniques.
@@ A.1 Additional Details for Methodology
In this section, we present additional details for the Methodology, including an illustration of batch split, the way to construct the global optimal label mapping for low entropy attack, the overall poisoning & surrogate model update algorithm, as well as details of defense practices.
@@ A.1.2 Global Optiomal Label Mapping
A.1.3 Detailed Poisoning & Training Algorithm

### v3

@@ 4.3 TTA-Aware Attack Objective
Low Entropy Attack Objective: High entropy attack objective is particular effective against self-training based TTA methods because of high updating gradient, yet they could be easily defended by some defense strategies, e.g. entropy thresholding. Therefore, we further explore a new low entropy based attack objective. DIA (Wu et al., 2023) proposed to maximize the cross-entropy loss on the benign samples (indiscriminate attack) to generate the other poisoned samples. However, we empirically found that maximizing the cross-entropy loss without any constraints on one sample is prone to maximizing the probability of the most confident class (except the ground-truth) of one model, and feeding these samples into TTA model would bring up the following issues. i) The model will quickly bias towards the most confident class and collapse if without any class diversity constraints. ii) If class diversity constraints are applied (assembled in several TTA methods, e.g. EATA, ROID), this objective will become less ineffective, since class-biased poisoned samples will obtain a less updating weighting than other benign samples. Therefore, we propose a class-balanced low entropy attack, termed Balanced Low Entropy Attack (BLE Attack). Specifically, we maintain an moving average probability confusion $C\in[0,1]^{K\times K}$ to store the prediction bias in each class and find a global optimal label mapping $M\in\{0,1\}^{K\times K}$ such that each class is attacked towards the most probable non ground-truth class. Details of deriving label mapping $M$ is deferred to the Appendix A.1.2. Finally, the BLE Attack objective is calculated as Eq. 8.
@@
Overall Attack Strategy: We craft poisoned data by attacking the aforementioned in-distribution attack objective with regularization. Following the practice that poisoned data should be less discernible by human, we employ a 40 steps Projected Gradient Descent algorithm (Boyd & Vandenberghe, 2004) on the combined objective in Eq. 6 with a budget $b$. More details of the whole data poisoning algorithm are deferred to the Appendix A.1.6.
@@ 5.1 Experiment Details
Benchmark Poisoning Methods: We evaluated the following methods under our proposed RTTDP setting. Unlearnable Examples (Huang et al., 2021) generates the poisoned noise by minimizing the cross-entropy of $\mathcal{B}_{a}$ on a randomly initialized model. Adversarial Poisoning (Fowl et al., 2021) proposed to minimize the cross-entropy between the posterior probabilities of poisoned samples, $\mathcal{B}_{a}$, and their corresponding incorrect labels, $\hat{y}$, where the incorrect labels are defined as $\hat{y}_{i}=y_{i}+1$. DIA (Wu et al., 2023) is one of the first approaches towards test-time data poisoning, generating the poisoned data via maximizing the cross-entropy of other benign data. We adapt DIA to the realistic evaluation protocol by splitting $\mathcal{B}_{ab}$ into two subsets of $50\%$ each, i.e. $\mathcal{B}^{p}_{ab}:\mathcal{B}^{b}_{ab}=1:1$, and craft $\mathcal{B}^{p}_{a}$ to maximize the cross-entropy of $\mathcal{B}^{b}_{ab}$. TePA (Cong et al., 2023) proposed to maximize entropy to generate poisoned data and performed attack before TTA starts. We adapt TePA to generate the poisoned data based on the source model and inject them into TTA pipeline on-the-fly. MaxCE (Madry et al., 2018) is an established way to create adversarial samples by maximizing the cross-entropy loss. Finally, we evaluate the two attack objectives proposed in this paper, i.e high entropy attack (NHE Attack) and low entropy attack against most probable and balanced non ground-truth class (BLE Attack). For both NHE Attack and BLE Attack, we evaluate the attack objective subject to the constraint of our proposed feature consistency (Eq. 6). For all methods that require gradient-based optimization, we employ the Projected Gradient Descent (PGD) algorithm (Boyd & Vandenberghe, 2004) to perform the constrained optimization. We use 40 steps PGD for all methods for a fair comparison.
Datasets: We evaluate on three datasets, widely adopted for TTA benchmarking. CIFAR10-C, CIFAR100-C and ImageNet-C are synthesized from the original clean validation set by adding various types of corruptions to simulate natural distribution shifts (Hendrycks & Dietterich, 2019). We choose corruption level 5 and perform continual test-time adaptation setting (Wang et al., 2022) for evaluation. Following prior works (Wang et al., 2022; Döbler et al., 2023), we adopt the pre-trained WideResNet-28 (Zagoruyko & Komodakis, 2016), ResNeXt-29 (Xie et al., 2017), and ResNet-50 (He et al., 2016) models for experiments on the CIFAR10-C, CIFAR100-C, and ImageNet-C datasets, respectively. More experiment details can be found in the Appendix A.3.
@@ 5.2 Evaluation on Test-Time Data Poisoning
Table 2: Evaluation of test-time data poisoning under the RTTDP protocol for CIFAR10-C. We report the attack success rate (higher the better) for each TTA method and the average ranking (lower the better) for each attack objective. indicates that the method is modified to align with the RTTDP protocol.
@@
Uniform No Attack 43.81 19.72 21.00 18.03 17.41 18.94 16.46 16.37 18.28 7.43
Unlearnable Examples (Huang et al., 2021) 32.61 26.62 20.11 18.43 19.23 17.27 17.80 21.72 4.86
Adversarial Poisoning (Fowl et al., 2021) 19.60 19.90 18.94 18.69 19.90 18.34 19.12 19.21 3.86
DIA (Wu et al., 2023) 26.04 21.87 18.94 18.56 19.46 17.72 17.77 20.05 4.86
TePA (Cong et al., 2023) 33.78 22.36 23.37 17.75 19.53 16.57 18.76 21.73 4.43
MaxCE (Madry et al., 2018) 18.55 20.81 18.17 18.50 19.50 16.88 18.57 18.71 6.00
BLE Attack (Ours) 54.07 51.99 45.20 34.00 26.80 18.12 19.06 35.61 1.57
NHE Attack (Ours) 73.86 72.40 29.73 18.67 24.56 17.54 17.00 36.25 2.86
@@
Table 3: Evaluation of test-time data poisoning under the RTTDP protocol for CIFAR100-C. We report the attack success rate (higher the better) for each TTA method and the average ranking (lower the better) for each attack objective. indicates that the method is modified to align with the RTTDP protocol.
@@
Uniform No Attack 46.23 60.25 47.12 32.20 31.93 31.62 32.13 29.11 37.77 7.29
Unlearnable Examples (Huang et al., 2021) 75.31 72.74 37.59 33.66 41.75 32.66 31.31 46.43 3.71
Adversarial Poisoning (Fowl et al., 2021) 33.87 34.27 32.11 35.49 32.27 32.76 30.15 32.99 5.71
DIA (Wu et al., 2023) 74.79 68.42 33.77 32.57 33.20 32.68 30.10 43.65 5.29
TePA (Cong et al., 2023) 75.79 81.98 36.74 33.41 35.79 32.41 32.24 46.91 3.57
MaxCE (Madry et al., 2018) 42.36 39.04 34.08 40.73 32.01 32.18 31.08 35.93 5.57
BLE Attack (Ours) 73.93 76.71 47.30 34.58 43.25 32.81 32.27 48.69 2.28
NHE Attack (Ours) 92.08 91.72 37.86 33.85 56.09 32.50 31.48 53.65 2.57
@@
Table 4: Evaluation of test-time data poisoning under the RTTDP protocol for ImageNet-C. We report the attack success rate (higher the better) for each TTA method and the average ranking (lower the better) for each attack objective. indicates that the method is modified to align with the RTTDP protocol.
@@ 5.3 Ablation Study on Attack Modules
In this section, we ablate our proposed modules including the surrogate model, the feature consistency regularization and two attack objectives to demonstrate their their indispensable contribution to the final results. We conduct the experiments on CIFAR10-C and CIFAR100-C datasets as shown in Tab. 5. First, comparing the use of the source model v.s. surrogate model for generating poisoned data, the surrogate model consistently delivers superior results, often approaching or even slightly surpassing those obtained with the online model, regardless of the attack objective. It demonstrates the effectiveness of our proposed surrogate model that is leveraged for generating on-the-fly poisoned data. Second, our proposed NHE attack objective is effective though using source model and without feature consistency regularization, that could be attributed to the high entropy samples easily mislead the model update and cause strong perturbation to the source knowledge. Third, under the surrogate model or online model, both BLE and NHE are significantly improved with the help of feature consistency regularization, empirically demonstrating the reasonableness and effectiveness of our method.
@@
TENT EATA SAR AR ($\uparrow$) TENT EATA SAR AR ($\uparrow$)
@@ 5.5 Generalisation of our modules
In this section, to demonstrate the generalization of our method, we combine the existing attack objectives with our proposed modules including the surrogate model and feature consistency regularization. The comparisons are shown in Tab. 7. First, we can observe that DIA and MaxCE could get improved additional with our proposed method. Second, TePA could obtain improvement under TENT method but slightly degraded under EATA method. It could be that TePA using maximizing entropy as an attack objective, and with the help of the surrogate model, the poisoned data would have very high prediction entropy because the attack reference model is more approximate to the online model, but fail to pass the entropy threshold and class diverse weighting using in EATA. Overall, our proposed in-distribution attack could generalize to most of the existing attacking objectives.
@@ 6 Conclusion
Acknowledgements
This research work is supported by the National Natural Science Foundation of China (NSFC) (Grant Number: 62106078), the Agency for Science, Technology and Research (A*STAR) under its MTC Programmatic Funds (Grant Number: M23L7b0021) and the Guangdong R&D key project of China (Grant Number: 2019B010155001). This work was done during Yongyi Su’s attachment with Institute for Infocomm Research (I2R), funded by China Scholarship Council (CSC).
@@ A.1.2 Global Optiomal Label Mapping
A.1.3 Details of Defense Practices
Here, we provide the details about the defense practices evaluated in Tab. 7 of the main text.
A.1.4 Distinction between RTTDP and TePA
We would like to highlight the key differences between our proposed RTTDP and TePA protocols as follows,
TePA (Cong et al., 2023) employs a fixed surrogate model before test-time adaptation begins for generating poisoning, which qualifies the method as an offline method. The surrogate model is obtained by training a separate model (different architecture from the target model) using the same source dataset. For example, on TTA for CIFAR10-C, if the target model, i.e. the model deployed for inference and is subject to test-time adaptation, is ResNet18, TePA employs VGG-11 as the surrogate model and trains VGG-11 on the same source training dataset (CIFAR10 clean training set). This is evidenced from the source code released by official repository  [^1] https://github.com/tianshuocong/TePA and the descriptions in TePA "we assume that the adversary has background knowledge of the distribution of the target model’s training dataset. This knowledge allows the adversary to construct a surrogate model with a similar distribution dataset".
TePA employs the fixed surrogate model to generate poisoned dataset $x^{\prime}$. Then generated poisoned dataset is fed to test-time adaptation to update model weights. Afterwards, TTA is further conducted on clean testing data for model update and performance evaluation. The segregation of data poisoning and TTA steps further support the claim that TePA should be classified as an offline approach.
Finally, to ensure fair comparison between TePA with our proposed methods under RTTDP protocol, TePA could be adapted to online fashion and we made such an adaptation to TePA for comparison in Tab. 2, Tab. 3 and Tab. 4 of the main text. Specifically, we use TePA to generate poisoning against the initial surrogate model and inject the generated poisoning into the testing data stream, i.e. placing poisoning in between benign testing batches. In this way, poisoning will affect TTA in an online fashion. We believe this is the most fair way to compare RTTDP with TePA.
A.1.5 More analysis and derivations about the optimization objective
Here, we discuss the transition from the original bi-level optimization objective (Eq. 2) to our proposed single-level optimization objective (Eq. 4) with a feature consistency constraint.
The original optimization objective for test-time data poisoning is formulated as a bi-level optimization problem, as shown below (equivalent to the meaning of Eq. 2):
[…]
The above optimization involves an inner loop where the model adapts to test samples, including poisoned and benign samples, and an outer loop to optimize the attack objective. This structure is computationally intensive and impractical under the constraints of the RTTDP setting. To address this, we provide a detailed step-by-step derivation and explanation below.
1. Discarding the Inner Optimization: In DIA (Wu et al., 2023), the inner optimization is approximated by assuming $\theta_{t}^{*}\approx\theta_{t}$, where $\theta_{t}^{*}$ represents the parameters after a full adaptation step, and $\theta_{t}$ represents the current parameters. This approximation is justified as TTA models typically update minimally during a single minibatch iteration, resulting in minor perturbations to $\theta_{t}$. Thus, the approximation retains practical relevance while simplifying the problem. The formula is derived as (this is also the DIA’s objective),
$\mathcal{B}_{a}=\arg\min_{\mathcal{B}_{a}}E_{(x,y)\in\mathcal{B}_{b}}\left[\mathcal{L}_{atk}(h(x;\theta_{t}(\mathcal{B}_{a}\cup\mathcal{B}_{b})),y)\right]$
2. Surrogate Model for Online Parameters: In the RTTDP protocol, direct access to online model parameters $\theta_{t}$ is unrealistic. Instead, we replace $\theta_{t}$ with the surrogate model parameters $\hat{\theta}_{t}$, which are accessible and trained to approximate the online model’s behavior.
$\mathcal{B}_{a}=\arg\min_{\mathcal{B}_{a}}E_{(x,y)\in\mathcal{B}_{b}}\left[\mathcal{L}_{atk}(h(x;\hat{\theta}_{t}(\mathcal{B}_{a}\cup\mathcal{B}_{b})),y)\right]$
3. Removing the access to $\mathcal{B}_{b}$: In the RTTDP protocol, the adversary is prohibited from observing benign users’ samples when generating poisoned samples. Consequently, the $\mathcal{B}_{b}$ term is excluded from the optimization objective. In the main text, we introduce to leverage $\mathcal{B}_{ab}$ to replace $\mathcal{B}_{b}$, where $\mathcal{B}_{ab}$ represents the adversary benign samples before they are poisoned.
$\mathcal{B}_{a}=\arg\min_{\mathcal{B}_{a}}E_{(x,y)\in\mathcal{B}_{ab}}\left[\mathcal{L}_{atk}(h(x;\hat{\theta}_{t}(\mathcal{B}_{a}\cup\mathcal{B}_{ab})),y)\right]$
where $\hat{\theta}_{t}(\mathcal{B}_{a}\cup\mathcal{B}_{b})$ indicates forwarding $\mathcal{B}_{a}\cup\mathcal{B}_{b}$ to update the BN statistics. This objective would lead to a trivial solution that $\mathcal{B}_{a}$ is effective only for the current $\mathcal{B}_{ab}$ data through easily introducing biased normalization in each BN layer, and it has little effect while $\mathcal{B}_{a}$ and $\mathcal{B}_{ab}$ are in seperated batch. Therefore, it would waste a half of attack query budget for forwarding these poisoned samples (the benign samples take up half of the batch size).
4. Introducing a feature consistency constraint to improve query utilization: In the main text, we observed the feature distributions of $\mathcal{B}_{a}$ and $\mathcal{B}_{ab}$ and found out that they obviously do not overlap, so we introduced feature consistency constraint to regularize their distributions according to the PAC learning framework in order to merge the two subsets into a single one, and to improve the utilization of the poisoned data query. The final objective is derived as follows, where $P_{a}$ and $P_{ab}$ are the shallow feature distributions of $\mathcal{B}_{a}$ and $\mathcal{B}_{ab}$, respectively.
$\mathcal{B}_{a}=\arg\min_{\mathcal{B}_{a}}E_{(x,y)\in\mathcal{B}_{a}}\left[\mathcal{L}_{atk}(h(x;\hat{\theta}_{t}(\mathcal{B}_{a})),y)\right],\ \ s.t.\ \ D(P_{a},P_{ab})=0$
To this end, we can fully utilize the budget of all poisoned data query to generate poisoned data. The experimental results show that the attack performance of the attack objective will be significantly improved after using this regularization term.
In-distribution Attack Objective from a TTA Perspective: Our proposed objective leverages the dependence of TTA models on self-training mechanisms, which aim to maximize confidence on pseudo-labels for adaptation. When the TTA model adapts to poisoned samples, it learns and reinforces incorrect associations. This creates a vulnerability, as future test samples with similar shallow feature distributions are more likely to be misclassified by the online model. Since TTA methods iteratively adapt using incoming test samples, our approach leverages this dependency to propagate the error induced by poisoned samples throughout the adaptation process.
A.1.6 Detailed Poisoning & Training Algorithm